# Preprint Manuscript

### v3

@@ 4.5 Proof Justification
This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is 11 and the rest are 0. ∎
@@ 4.6 Target Formula Constraint
Let $G_{T}$ be the target formula with Gödel number $g_{T}$. Enforce, with $T$ is the global multiplier (Definition 3.8),
$T\cdot(f_{N}-g_{T})$ $=0,$ (4.72)
$T-1-U^{2}$ $=0,$ (4.73)
@@
Constraint 4.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
Proof.
Since $T=1+U^{2}\geq 1$ (by equation (4.73)), equation (4.72) forces $f_{N}=g_{T}$. ∎
@@ 4.7 Complete System Assembly
Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{T}$ be a sentence. There exists a finite system $\Phi(\mathbf{x})$ of polynomial equations over $\mathbb{N}$, each of degree at most $3$, such that:
$T_{\text{Fib}}\vdash G_{T}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (4.75)
Proof.
@@
    Suppose $T_{\text{Fib}}\vdash G_{T}$. Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{T}$. Let $f_{i}$ be the Gödel number of $\varphi_{i}$. By Theorem 3.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that $\beta(c^{*},d^{*},i)=f_{i}$ for all $i\leq N$. For each $f_{i}$, compute its Zeckendorf representation $d^{*}_{i,\kappa}$ (which exists and is unique by Definition 3.3). For each line $i$:
@@
    Compute the Fibonacci addition witnesses (carries $c^{*}_{i,j,k,\kappa}$, sums $s^{*}_{i,j,k,\kappa}$) Finally, compute all guard variables $u^{*}_{\bullet},v^{*}_{\bullet}$ to satisfy definitions 3.63.8. By construction, $\mathbf{x}^{*}=(N,c^{*},d^{*},\{f^{*}_{i}\},\ldots)$ satisfies
    $\Phi(\mathrm{x}^{*})=\mathbf{0}.$
@@
    Let $\varphi_{i}$ be the formula with Gödel number $f^{*}_{i}$. The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{T}$ in $T_{\text{Fib}}$, since:
@@
    Therefore, $T_{\text{Fib}}\vdash G_{T}$.
@@ 5.1 Merging via Sum-of-Squares
Degree bound. By Table 2, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.6).
Merger. The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
@@
Size. With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
@@ 5.2 Degree Reduction to Cubic
Proof sketch.
For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-$2$ constraint. Iterate until all monomials have degree $\leq 3$. See [Jones, 1982] for details. ∎
Applying Theorem 5.1 to $P_{\text{merged}}$ (degree $6$, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
@@ 5.3 Main Result
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{T}$ its Gödel sentence. There exists a polynomial
@@
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (5.78)
@@
Remark (Non-Uniformity).
The polynomial $P$ depends on the (unknown) proof length $N$. This is inherent: no algorithm can compute the “correct” $P$ from $T$ and $G_{T}$ alone, as this would solve the Halting Problem.
@@ 6.1 Degree Reduction Protocol
Then $\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ and $Q_{E}=0\iff E=0$ when $U\geq 1$.
@@ 6.3 Solvability and Complexity
Remark (Application to $P_{\text{merged}}$).
For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 2), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums. Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(K^{3}N^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
@@ 7 Result
We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 4 faithfully encodes provability in $T_{\text{Fib}}$. The proof proceeds bidirectionally through soundness (Section 7.1) and completeness (Section 7.2), yielding the main result (Theorem 5.2).
@@ 7.1 Soundness
If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $T_{\text{Fib}}\vdash G_{T}$.
Proof.
@@
Step 3. Zeckendorf representability. By Lemma 4.3, each $f_{i}^{*}$ admits a unique representation:
@@
with Zeckendorf digits computed via Exposition Exposition.
@@
$f_{N^{*}}^{*}=g_{T}.$ (7.86)
Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{T}$ in $T_{\text{Fib}}$. ∎
@@ 7.2 Completeness
If $\mathrm{T}_{\text{Fib}}\vdash G_{T}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
Proof.
Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{T}$. We construct $\vec{x}^{*}$ by explicit assignment.
@@
    Lemma 4.3 holds (via Zeckendorf uniqueness, Definition 3.3),
@@
    Lemma 4.13 holds (since $f_{N}^{*}=g_{T}$).
@@
7.3 Main Algorithm: Encoding Gödel’s Theorem
@@
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.90)
Proof.
@@
7.4 Undecidability Consequences
@@
Proof.
By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 5.2), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability. ∎
@@
There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing machine $M$ with Gödel number $\ulcorner M\urcorner$:
$\ltxml@oqmark@open\textquotedblleft\penalty 10000\thinspace\thinspace M\text{ halts on empty input}\textquotedblright\ltxml@oqmark@close{}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (7.91)
@@
Proof sketch.
By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing machine halting. Applying the guard-gadget framework (Section 3.5) and degree reduction (Theorem 5.1) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability. The map $M\mapsto\Phi(\ulcorner M\urcorner)$ is computable by effective construction of guards and auxiliary variables. ∎
7.5 Impredicativity
Thesis 7.6 (Impredicativity of Degree Threshold).
@@ 7.5 Impredicativity
Thesis 7.7 (Diophantine Impredicativity).
@@
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 7.6, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
    Externalize the threshold: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel sentences not formalizable in $\mathsf{Q}$), or
@@
    Adopt impredicative axioms: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
@@
    No uniform refutation exists (by Thesis 7.6, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
Thus, the threshold constitutes a Brouwerian counterexample to the law of excluded middle: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof. The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the halting problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
Thesis 7.9 (Impredicativity of Degree Threshold).
@@
This result formalizes the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative. Any attempt to define the “first” degree $k$ at which undecidability arises leads to a form of self-reference: specifying such a $k$ enables the construction of a Diophantine equation of degree $k$ whose solvability encodes the negation of that very definition. Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory. This mirrors classical semantic paradoxes—such as the Grelling–Nelson paradox [^6] Consult [Quine, 1976] for a thorough discussion of paradoxes.—in which the act of defining:
Is the adjective ’heterological’ a heterological word?
yields contradiction. Similarly, “Does undecidability start at kk?” becomes undecidable when answered affirmatively. If $k$ is the first undecidable degree, the evaluation criterion shifts, making the boundary undecidable if and only if it is decidable—a fixed point in arithmetic. Under HBK [^7] The Heyting–Brouwer–Kolmogorov (HBK), formalized by Heyting [1930], omits proof by the law of excluded middle unless constructively justified., this reflects the failure by appealing to the law of excluded middle (LEM): asserting $\mathsf{Dec}(k)\vee\neg\mathsf{Dec}(k)$ requires a uniform method for deciding solvability across all equations in $\mathcal{D}_{k}$, or uniformly refuting such a method—a totality that can be rejected without explicit witnesses.
Thus, impredicativity arises not merely from syntactic self-reference, but from the semantic attempt to define a global threshold for undecidability. The paradox is resolved constructively by recognizing that no such threshold is constructively definable: the transition from decidable to undecidable is not marked by a computable boundary, but by a failure of totality. In this sense, the undecidability threshold is a meta-mathematical horizon—a boundary that cannot be crossed without stepping outside the constructive framework. Any attempt to define it collapses into a fixed point of the form:

### v4

@@ 4.5 Proof Justification
This is a linear Diophantine equation with boolean variables, which has a solution if and only if exactly one variable is $1$ and the rest are 0. ∎
@@ 4.6 Target Formula Constraint
Let $G_{\mathrm{T}}$ be the target formula with Gödel Number $g_{\mathrm{T}}$. Enforce, with $\mathrm{T}$ is the global multiplier (Definition 3.11),
$\mathrm{T}\cdot(f_{N}-g_{\mathrm{T}})$ $=0,$ (4.92)
$\mathrm{T}-1-U^{2}$ $=0,$ (4.93)
@@
Constraint 4.13 forces $f_{N}=g_{\mathrm{T}}$ (i.e., the final line of the proof is the target formula).
Since $\mathrm{T}=1+U^{2}\geq 1$ (by equation (4.93)), equation (4.92) forces $f_{N}=g_{\mathrm{T}}$. ∎
@@ 4.7 Complete System Assembly
Let $T_{\text{Fib}}$ be a recursively axiomatizable theory with finite axiom set $\mathcal{A}$, and let $G_{\mathrm{T}}$ be a sentence. There exists a finite system $\Phi(\mathbf{x})$ of polynomial equations over $\mathbb{N}$, each of degree at most $3$, such that:
$T_{\text{Fib}}\vdash G_{\mathrm{T}}\quad\iff\quad\exists\mathbf{x}\in\mathbb{N}^{m}:\Phi(\mathbf{x})=\mathbf{0}.$ (4.95)
@@
    Suppose $T_{\text{Fib}}\vdash G_{\mathrm{T}}$. Then there exists a finite proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ where $\varphi_{N}=G_{\mathrm{T}}$. Let $f_{i}$ be the Gödel Number of $\varphi_{i}$. By Theorem 3.1, there exist $c^{*},d^{*}\in\mathbb{N}$ such that $\beta(c^{*},d^{*},i)=f_{i}$ for all $i\leq N$. For each $f_{i}$, compute its Zeckendorf Representation $d^{*}_{i,\kappa}$ (which exists and is unique by Definition 3.3). For each line $i$:
@@
    For each MP line $i$ with premises $(j,k)$, compute the Zeckendorf Digits of the value $f_{j}+f_{k}$ using Procedure 4.9 and assign $d_{i,\kappa}^{*}$ accordingly. No auxiliary carry or sum variables are introduced: the constraints enforce $f_{i}=f_{j}+f_{k}$ and then verify that $(d_{i,\kappa}^{*})$ is the unique Zeckendorf vector for that value (Exposition Exposition).
@@
    Let $\varphi_{i}$ be the formula with Gödel Number $f^{*}_{i}$. The sequence $\langle\varphi_{1},\ldots,\varphi_{N^{*}}\rangle$ is a valid proof of $G_{\mathrm{T}}$ in $\mathrm{T}_{\text{Fib}}$, since:
@@
    Therefore, $T_{\text{Fib}}\vdash G_{\mathrm{T}}$.
@@ 5.1 Merging via Sum-of-Squares
Degree bound: By Table 1, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.9). Merger: The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
@@
Size: With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
@@ 5.2 Degree Reduction to Cubic
For each monomial $\mathbf{x}^{\mathbf{a}}$ with $|\mathbf{a}|>3$, introduce $y$ via $y=\prod_{j\in S}x_{j}^{a_{j}}$ (where $|\mathbf{a}|_{S}\geq 2$), expressed as a degree-22 constraint. Iterate until all monomials have degree $\leq 3$. See [Jones, 1982] for details. ∎ Applying Theorem 5.2 to $P_{\text{merged}}$ (degree 66, $O(KN^{9})$ monomials by squaring analysis) introduces $k=O(KN^{9})$ auxiliary variables.
@@ 5.3 Main Result
Let $\mathrm{T}$ be a recursively axiomatizable theory with finite axioms, and $G_{\mathrm{T}}$ its Gödel Sentence. There exists a polynomial
@@
$\mathrm{T}\vdash G_{\mathrm{T}}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P(\mathbf{x})=0.$ (5.98)
@@
Remark.
The polynomial $P$ depends on the (unknown) proof length $N$. This is inherent: no algorithm can compute the “correct” $P$ from $\mathrm{T}$ and $G_{\mathrm{T}}$ alone, as this would solve the Halting Problem.
@@ 6.1 Degree Reduction Protocol
Then
$\delta(Q_{E})=\max(1+\delta(E),2)\leq 3$ (6.100)
and
$Q_{E}=0\iff E=0$ (6.101)
when $U\geq 1$.
@@ 6.2 Solvability and Complexity
Remark.
For
$P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ (6.104)
with $\delta(P_{i})\leq 3$ (from Table 1), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore
$m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2}),$ (6.105)
where
$\bar{s}=\max_{i}s_{i}=O(K)$ (6.106)
by the bounds on Zeckendorf Digit sums. Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(K^{3}N^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
@@ 7 Result
We establish the central equivalence: the polynomial $P_{\text{hard}}$ constructed in Section 4 faithfully encodes provability in $\mathrm{T}_{\text{Fib}}$. The proof proceeds bidirectionally through soundness (Section 7.1) and completeness (Section 7.2), yielding the main result (Theorem 5.3).
@@ 7.1 Soundness
If there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$, then $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$.
@@
Step 3. Representability. By Lemma 4.3, each $f_{i}^{*}$ admits a unique representation:
@@
$f_{N^{*}}^{*}=g_{\mathrm{T}}.$ (7.111)
Conclusion. The sequence $\langle f_{1}^{*},\ldots,f_{N^{*}}^{*}\rangle$ constitutes a valid proof of $G_{\mathrm{T}}$ in $T_{\text{Fib}}$. ∎
@@ 7.2 Completeness
If $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$, then there exists $\vec{x}^{*}\in\mathbb{N}^{M}$ such that $P_{\text{hard}}(\vec{x}^{*})=0$.
Let $\pi=\langle f_{1},\ldots,f_{N}\rangle$ be a proof of $G_{\mathrm{T}}$. We construct $\vec{x}^{*}$ by explicit assignment.
@@
    Lemma 4.3 holds (via uniqueness, Definition 3.3),
@@
    Lemma 4.13 holds (since $f_{N}^{*}=g_{\mathrm{T}}$).
@@
7.3 Encoding Incompleteness
@@
$\mathrm{T}\vdash G_{\mathrm{T}}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.115)
@@
7.4 Calibration
Logic 7.4.
It is useful to record a simple calibration: a hypothetical total correct solver for the cubic fragment would already decide solvability for degree $\leq 4$ equations by a uniform compilation. Concretely, there is a computable, uniform transformation which maps any quartic equation
$P(\vec{x})=0\qquad(\delta(P)\leq 4)$ (7.116)
to a cubic equation
$Q(\vec{x},\vec{y})=0\qquad(\delta(Q)\leq 3)$ (7.117)
such that
$\exists\vec{x}\in\mathbb{N}^{n}\;P(\vec{x})=0\quad\Longleftrightarrow\quad\exists(\vec{x},\vec{y})\in\mathbb{N}^{n+m}\;Q(\vec{x},\vec{y})=0.$ (7.118)
The transformation is the standard shielding discipline: each quartic monomial is replaced by a fresh auxiliary variable, constrained by a short chain of quadratic/biquadratic equations (introducing only bounded-degree constraints). One then packages the resulting finite system into a single Diophantine equation (e.g. by a sum-of-squares merger), and applies the same degree-reduction protocol already used in our aggregation layer to return to total degree $\leq 3$.
Thus, if $\exists e\,\mathsf{Solve}_{3}(e)$ held, then degree-$4$ solvability over $\mathbb{N}$ would be decidable as an immediate consequence. This is not our theorem, but it underscores the strength of the cubic solver hypothesis: it would already collapse the known degree-$4$ regime into the cubic one by a uniform compilation.
Logic 7.5.
By contrast, there is no uniform compilation from the quadratic to the cubic fragment. Quadratic Diophantine equations are closed under addition and squaring but not under multiplicative composition. Any attempt to encode a cubic monomial $x_{i}x_{j}x_{k}$ using only quadratic constraints (for example $y=x_{i}x_{j}$, $z=yx_{k}$) inevitably introduces a constraint of degree $3$. Thus multiplication itself is not quadratic-definable.
Formally, the quadratic fragment defines only semilinear subsets of $\mathbb{N}^{n}$, while the cubic fragment already defines recursively enumerable sets. The gap is therefore categorical rather than technical: degree $3$ is the first level at which bounded verification becomes representable, and where existential closure fails uniformly.
7.5 Cubic Incompleteness
@@
By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{\mathrm{T}}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{\mathrm{T}}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 5.3), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Halting Problems via similar constructions yields a uniform reduction from the Halting Problem to $\mathcal{D}_{3}$ solvability. ∎
@@
There exists a computable function $\Phi:\mathbb{N}\to\mathbb{Z}[\vec{x}]$ such that for any Turing Machine, cf. Turing [1937, 1938] $M$ with Gödel Number $\ulcorner M\urcorner$:
$M\text{ halts on empty input}\quad\Longleftrightarrow\quad\exists\vec{x}\in\mathbb{Z}^{k}:\Phi(\ulcorner M\urcorner)(\vec{x})=0,$ (7.119)
@@
By [Matiyasevich, 1970], there exists a polynomial $P_{\text{MRDP}}(m,\vec{y})$ encoding Turing Machine. Applying the guard-gadget framework and degree reduction (Theorem 5.2) yields a cubic polynomial $\Phi(\ulcorner M\urcorner)$ preserving solvability. The map $M\mapsto\Phi(\ulcorner M\urcorner)$ is computable by effective construction of guards and auxiliary variables. ∎
Thesis 7.8 (Impredicativity of Degree Threshold).
@@
Lemma 7.9 (Compactness–Reflection Barrier).
Let
$\Phi_{e,p,t,b}(\vec{z})=0$ (7.121)
be the cubic system from Theorem 2.7 such that
$\exists\vec{z}\;\Phi_{e,p,t,b}(\vec{z})=0\Longleftrightarrow\mathsf{Run}(e,p,t,b).$ (7.122)
Fix $(e,p,b)$ and write
$\mathsf{HaltsOut}(e,p,b)\equiv\exists t\,\mathsf{Run}(e,p,t,b)$ (7.123)
  1. (i)
    Bounded coherence. For each bound $B\in\mathbb{N}$, the bounded predicate
    $\mathsf{HaltsOut}_{\leq B}(e,p,b)\;:\Longleftrightarrow\;\exists t\leq B\;\mathsf{Run}(e,p,t,b)$ (7.124)
    is decidable, and is representable by the finite disjunction
    $\mathsf{HaltsOut}_{\leq B}(e,p,b)\;\Longleftrightarrow\;\bigvee_{t\leq B}\;\exists\vec{z}\;\Phi_{e,p,t,b}(\vec{z})=0.$ (7.125)
  2. (ii)
    Compactness of the limit step. In any classical metatheory with $\omega$-completeness/compactness for $\Sigma_{1}$-instances (equivalently: a principle validating passage from all bounded stages to the unbounded existential),
    $\Bigl(\forall B\;\mathsf{HaltsOut}_{\leq B}(e,p,b)\Bigr)\ \Longrightarrow\ \mathsf{HaltsOut}(e,p,b).$ (7.126)
    Thus the unbounded existential closure is exactly the limit of its bounded approximants.
  3. (iii)
    Reflection constraint. Any argument that derives the global negation
    $\neg\exists\mathsf{Close}\;\forall e,p,b\;\bigl(\mathsf{HaltsOut}(e,p,b)\leftrightarrow\mathsf{Sol}_{3}(\mathsf{Close}(e,p,b))\bigr)$ (7.127)
    from bounded information alone must, at some point, employ the limit step of (ii) (or an equivalent reflection/compactness principle). In particular, the transition from the family
    $\{\mathsf{HaltsOut}_{\leq B}\}_{B\in\mathbb{N}}$ (7.128)
    to the global $\mathsf{HaltsOut}$ reintroduces precisely the unbounded closure move the cubic analysis isolates.
Corollary 7.10 (Reflexivity of the Cubic Barrier).
The obstruction is self-similar: proving the impossibility of a uniform cubic closure operator requires a metatheoretic compact Reflection Principle that performs the same global unbounding step that a putative uniform solver would have to internalize.
@@
Thesis 7.11 (Diophantine Impredicativity).
@@
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 7.8, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
    “Externalize the threshold”: Define $B(k)$ in a strictly stronger metatheory (e.g., $\mathsf{PA}$ proves undecidability of $\mathcal{D}_{3}$ using Gödel Sentences not formalizable in $\mathsf{Q}$), or
@@
    “Adopt impredicative axioms”: E.g., $\mathsf{ZFC}$ proves existence of a least undecidable degree by quantifying over all sets of equations via the power set axiom (which is impredicative).
@@
    No uniform refutation exists (by Thesis 7.8, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
Thus, the threshold constitutes a “Brouwerian” counterexample to the Law of Excluded Middle, cf. Brouwer [1908]; van Atten and Sundholm [2017]: a statement $P$ such that neither $P$ nor $\neg P$ admits constructive proof. The degree-$3$ bound is visible only from a classical metatheory (e.g., $\mathsf{ZFC}$ proves $\mathcal{D}_{3}$ is undecidable” by encoding the Halting Problem), but no predicative subsystem of $\mathsf{PA}$ can internalize this fact.
Thesis 7.13 (Impredicativity of Degree Threshold).
@@
This result formalizes the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative. Any attempt to define the “first” degree $k$ at which undecidability arises leads to a form of self-reference: specifying such a $k$ enables the construction of a Diophantine equation of degree $k$ whose solvability encodes the negation of that very definition. Consequently, the threshold of undecidability is not first-order definable within any sufficiently strong theory, as it is the act of defining: